OpenStudy (bananas1234):

Solve 4r^2 – 9 = 91. {± 3} {± 4} {± 5} {±6}

rvc (rvc):

\[4r^2-9=91\]

rvc (rvc):

bring the -9 to the RHS

OpenStudy (imtiaz7):

4x^2-9=91 4x^2=91+9 4x^2=100 x^2=100/4 x^2=25 x=_+5

OpenStudy (imtiaz7):

so c is your right answer

OpenStudy (bananas1234):

i think that makes sence, thank you

OpenStudy (imtiaz7):

welcome bananas :) can you please send me a milk shake of banana bc here weather is too hot lol

OpenStudy (imtiaz7):

so where are you am from Asia lol

OpenStudy (bananas1234):

im from somewhere :)

OpenStudy (imtiaz7):

okay then somewhere..

OpenStudy (bananas1234):

got to get back to school work

rvc (rvc):

well u are not supposed to give direct answers :)

OpenStudy (bananas1234):

yes.